An unknown amount of copper (c= 0.3851 J/g°C) is initially heated to 100.0 °C. It is dropped into a calorimeter (c = 1.40 J/g°C, m = 5.0 g) that is initially at 25.0°C. The final temperature of the experiment is 27.5°C. The mass of the water in the calorimeter is 2.0 x 10^2 g. How much does the sample of copper weigh?
1
Expert's answer
2019-10-15T08:36:09-0400

Solution.

"\theta \ is final \ temperature."

t1 = t2 = 25, because, initially, the calorimeter and the water in it were in equilibrium, therefore their initial temperatures are equal.

"Q1 = c(H2O) \times m(H2O) \times (\theta-t1)"

"Q2 = c(calor.) \times m(calor.) \times (\theta-t2)"

"Q3 = c(Cu) \times m(Cu) \times (t3-\theta)"

"Q1 + Q2 = Q3"

We express the mass of cupper from here:

"m(Cu) = \frac{(4.2 \times 2*10^2 \times 2.5) + (1.4 \times 5 \times 2.5)}{0.3851 \times 72.5} = 75.84 \ g"

Answer:

m(Cu) = 75.84 g